OpenStudy (anonymous):

evaluate the algebraic expression for the given value or values of the variable(s) 1+5(x-4)^3; x=6

hero (hero):

Plug in x = 6 into the equation, then simplify: 1 + 5(6 - 4)^3

OpenStudy (anonymous):

is it 48?

hero (hero):

1 + 5(6 - 4)^3 = 1 + 5(2)^3 = 1 + 5(8) = 1 + 40 = 41

OpenStudy (anonymous):

thanks

hero (hero):

yw